SOHO and LASCO/EIT Recovery

On June 24, 1998, SOHO had a series of events that caused it to lose Sun pointing on June 25. The spacecraft started spinning around the solar vector sufficiently fast that it flipped to place the angular momentum vector with the highest moment of inertia pointing towards the Sun. Unfortunately, this meant that the solar arrays (and instruments) were pointed almost 90 degrees away from the Sun so that no power was being generated! This lack of power combined with the pointing change meant that no communication with the spacecraft was possible.

For the next several months, all we could do was wait. The spacecraft orbits around the Sun just like Earth. Eventually, the spacecraft's solar arrays would be pointed at the Sun. Contact with the spacecraft was made on July 23, 1998 by bouncing a signal off the spacecraft by the Arecibo and NASA/Deep Space Network RADAR. This confirmed the orbital position and the spacecraft spin rate, giving a tremendous emotional lift to the recovery team after having absolutely no contact with the spacecraft for a month. This information confirmed the results of detailed analyses that the spacecraft was spinning at about 1 RPM and in such a way that the solar arrays would be pointed at the Sun for a maximum of 30 seconds each spin period.

Then on August 3, a very short carrier signal from the spacecraft was received by the DSN. and 5 days later a short burst of telemetry was received. The telemetry allowed us to see our temperatures. LASCO temperatures were indicating that the instrument was colder than -50C which was the lowest limit that could be read out by our thermometer, a thermistor. An analysis by the thermal engineers showed that the LASCO was actually between -80 to -120C. The CCD cameras were at the colder end of this range.

On September 16, 1998, SOHO reacquired the Sun in a very dramatic fashion. The time-line of the steps for the reacquisition procedure was put onto the web, allowing the whole world to feel the excitement. The steps were completed without a hitch. The ESTEC, MATRA, and NASA engineers did a great job. The LASCO optical and electronics boxes (COB and LEB) came out of the deep freeze that they were in and soaked at the nominal operating temperatures for several weeks.

We started the LASCO/EIT instrument recovery at about 1500 UT on October 13, 1998. The progress of the instrument recovery can be followed on the Reacquisition Status page and detailed status reports.

For more information, follow the links below.